OpenStudy (anonymous):

Roseanna did not pay last month's credit card bill in full. Her daily balances for her last billing cycle are listed. What was Roseanna's average daily balance to the nearest cent? (Hint: Find the average or mean of these numbers.) For seven days she owed $456.11 For three days, she owed $1,077.00 For six days, she owed $900.08 For nine days, she owed $2,113.15 For five days, she owed $2,310.15

OpenStudy (aaronandyson):

are there any options given?

OpenStudy (aaronandyson):

The answer is $1413.11 Because 7*456.11=3192.77 3*1077.00=3231 6*900.08=5400.48 9*2113.15=19018.35 5*2310.15=11550.75 add that all up 3192.77+3231+5400.48+19018.35+11550.75=4... and then divide 42393.35/30=1413.111667

OpenStudy (anonymous):

it says it wtong

OpenStudy (aaronandyson):

I can see why your first instinct would be to add up all of the numbers and then divide by the number of days, but that is apparently wrong. For the first seven days, her balance was approximately $65.16 per day. (456.11 / 7). Then, she owed $359 per day for the next 3 days ($1,077/3), $150.01 for the next 6 days, $234.79 for the next 9 days, and then $462.03 for the final 5 days. If I am correct so far, you would then add each of these numbers up (65.16+359+150.01+234.79+462.03) and then divide that total by 5, (the number of numbers you added together). Your final answer would be $254.20. That is if and only if I was correct. I haven't done a math problem like that in a long time! Hope that if this isn't correct that you find your answer soon!